# Length-Induced Embedding Collapse in Transformer-based Models

### v1

@@ 5 Related Work
Text Embedding Models. Embeddings are generated by pre-trained language models (PLMs), which produce fixed-dimensional embeddings regardless of the input text length, laying the foundation of numerous NLP applications. Early works on text embeddings lack context awareness and are thus commonly labeled as word embedding models (Pennington et al., 2014). Modern embedding models (Wang et al., 2022a; Xiao et al., 2023) incorporate context awareness into language models through self-attention mechanisms, serving as the foundation for the latest embedding models. These models, based on Transformer architectures like BERT (Kenton & Toutanova, 2019) and RoBERTa (Liu et al., 2019), are first pre-trained on large-scale weakly supervised text pairs using contrastive loss (Gao et al., 2021), then fine-tuned on small scale but high-quality datasets. More recently, (Muennighoff et al., 2024) investigates the integration of generative and embedding tasks in large language models, introducing transformer-based GritLM, which enhances performance in both areas. Among the diverse range of pre-trained transformers (Wolf et al., 2020), self-attention plays a crucial role, and this paper focuses on how this module contributes to length collapse.
@@
Context Window Extension for Embedding Model. Despite transformer-based embedding models excelling in generating vector representations, they are typically constrained by a narrow context window of around 512 input tokens (Wang et al., 2022a; Xiao et al., 2023; Ni et al., 2022). This limitation significantly restricts their use in scenarios that require processing long inputs, such as extensive Wikipedia entries or meeting transcripts (Saad-Falcon et al., 2024; Zhu et al., 2024). Previous work attributes this poor performance to limited context windows and attempts to extend the window size. Current initiatives to develop long-context embedding models generally begin with acquiring a long-context backbone model, either by pre-training from scratch with long inputs (Günther et al., 2023; Nussbaum et al., 2024; Chen et al., 2024) or utilizing existing models (Wang et al., 2024; Zhu et al., 2024). This is followed by training the backbone model to generate embeddings. However, in this paper, we discover that regardless of the model’s context window size, the model consistently performs worse on longer texts than on shorter texts due to length collapse. Therefore, we aim to improve the performance of long texts across all context window size models by analyzing and addressing length collapse, rather than simply expanding context window size.
6 Conclusion and Future Work
@@
In this paper, we identify the phenomenon of length collapse, where embeddings of longer texts tend to cluster together and propose practical solutions through Fourier domain analysis. Our theoretical findings suggest that Multi-Head Self-Attention inherently performs stronger low-pass filtering as the token sequence length increases, leading to patch uniformity issues in longer sentences. To this end, we propose a technique called TempScale, which effectively reduces the low-pass filtering effect by introducing a temperature parameter when applying softmax in the self-attention matrix. Our extensive experiments validate the effectiveness of our methods and enhance the performance of general embedding models on the MTEB and LongEmbed benchmarks. Overall, TempScale is a crucial advancement in enhancing the performance of embedding models on longer texts.
@@
Future work includes: 1) LLM-based embedding model: Current work focuses on embedding models with bidirectional attention mechanisms. In the future, we plan to investigate length collapse in LLM-based embedding models that utilize unidirectional attention mechanisms; 2) Tuning method: The work in this paper relies on existing models and pre-trained parameters without using a training dataset. In future work, we will focus on tuning the temperature for additional improvements. 3) Analysis on more modules: We primarily investigate the impact of the self-attention module on length collapse in this paper. Moving forward, we plan to explore the effects of additional modules in transformers such as LayerNorm and FFN.
@@ Appendix A Background Information about Fourier Analysis
$\bm{DFT}=\frac{1}{\sqrt{n}}\begin{bmatrix}1&1&\cdots&1\\ 1&e^{2\pi\mathrm{j}}&\cdots&e^{2\pi\mathrm{j}(n-1)}\\ \vdots&\vdots&\ddots&\vdots\\ 1&e^{2\pi\mathrm{j}(k-1)\cdot 1}&\cdots&e^{2\pi\mathrm{j}(k-1)\cdot(n-1)}\\ \vdots&\vdots&\ddots&\vdots\\ 1&e^{2\pi\mathrm{j}(n-1)}&\cdots&e^{2\pi\mathrm{j}(n-1)^{2}}\end{bmatrix},$
@@
Hereby, we can independently operators $\mathcal{DC}\left[{\cdot}\right]$ and $\mathcal{HC}\left[{\cdot}\right]$ on echo channel using the matrices in Eqn. A. Then we can write $\mathcal{DC}\left[{\cdot}\right]$ as below:
@@ B.1 Proof of Theorem 2
Denote $\bm{B}=\begin{pmatrix}\bm{b_{1}}&\cdots&\bm{b_{n}}\end{pmatrix}$ and we have $\bm{AB}=\begin{pmatrix}\bm{Ab_{1}}&\cdots&\bm{Ab_{n}}\end{pmatrix}$. From the definition of the spectral norm, we have:$\lVert\bm{A}\rVert_{2}\geq\frac{\lVert\bm{Ab_{i}}\rVert_{2}}{\lVert\bm{b_{i}}\rVert_{2}}.$ Taking the average of the right-hand side, we obtain: $\lVert\bm{A}\rVert_{2}^{2}\geq\sum_{i=1}^{n}\frac{\lVert\bm{b_{i}}\rVert_{2}^{2}}{\lVert\bm{B}\rVert_{F}^{2}}\frac{\lVert\bm{Ab_{i}}\rVert_{2}^{2}}{\lVert\bm{b_{i}}\rVert_{2}^{2}}.$ This implies: $\lVert\bm{A}\rVert_{2}^{2}\lVert\bm{B}\rVert_{F}^{2}\geq\sum_{i=1}^{n}\lVert\bm{Ab_{i}}\rVert_{2}^{2}=\lVert\bm{AB}\rVert_{F}^{2}.$ Finally, the last step utilizes the result $\lVert\bm{A}\rVert^{2}_{F}=\sum_{i,j}|a_{ij}|^{2}=\sum_{j=1}^{n}\lVert\bm{a_{j}}\rVert^{2}_{2}$. Because both the spectral norm and the Frobenius norm of a matrix remain unchanged under transposition, we have $\lVert\bm{AB}\rVert_{F}=\lVert\bm{B^{\top}A^{\top}}\rVert_{F}\leq\lVert\bm{B^{T}}\rVert_{2}\lVert\bm{A^{T}}\rVert_{F}=\lVert\bm{A}\rVert_{F}\lVert\bm{B}\rVert_{2}$.
@@
$=\frac{1}{n}\left(\bm{I}-\frac{1}{n}\bm{11}^{T}\right)\bm{A11}^{\top}\bm{X}\bm{W}_{V}+\left(\bm{I}-\frac{1}{n}\bm{11}^{T}\right)\bm{AHW}_{V}$ (10)
@@ B.2 Proof of Theorem 3
(Filter Rate of Different Input Length $n$) Let $\bm{XW}_{Q}$ and $\bm{XW}_{K}$ be a Gaussian matrix, where elements $q_{ij}\sim\mathcal{N}(0,\sigma_{q}^{2})$ and $k_{ij}\sim\mathcal{N}(0,\sigma_{k}^{2}),\forall i,j.$ Let $x_{ij}=\bm{q}_{i}^{\top}\bm{k}_{j}/\sqrt{d}$ the attention score of pair $i,j$, whose variance can be expressed as $\sigma_{s}^{2}=\sigma_{q}^{2}\sigma_{k}^{2}+C_{cross}$, where $C_{cross}$ is the cross-covariance of the squared queries and keys (Goodman, 1960). Then we have
@@
The Eqn.  18 leverages $\left\lVert\bm{I}-\frac{1}{n}\bm{11}^{T}\right\rVert_{2}=1$ and $\lVert\bm{A}\rVert_{2}\leq\lVert\bm{A}\rVert_{F}$. Now we need to upper bound $\lVert\bm{A}\rVert_{F}$. Generally, the product of two independent Gaussian variables has a density in the form of a modiffed Bessel function of the second kind (Nahshan et al., 2024). When the vector dimensions are sufficiently large, the Central Limit Theorem implies that the distribution of the dot product between $\bm{q}_{i}$ and $\bm{k}_{j}$ can be approximated by a Gaussian distribution with zero mean and variance $\sigma^{2}_{s}$. As mentioned in Theorem 3, the variance of $\bm{q}_{i}^{\top}\bm{k}_{j}$ can be expressed as $\sigma^{2}=\sigma^{2}_{q}\sigma^{2}_{k}+C_{\rm cross}$, where $C_{\rm cross}=\text{Cov}(\bm{q}^{2},\bm{k}^{2})-\text{Cov}(\bm{q},\bm{k})^{2}$ is the cross-covariance of the squared queries and keys (Goodman, 1960). Thus we can suppose that each element $p_{j}\sim\mathcal{N}(0,\sigma_{s})$ in the matrix $\bm{XW_{Q}W_{K}^{\top}X^{T}}$ is independent, where $j\in(1,\cdots,n)$:
@@
$\sigma_{\Sigma}^{2}=\ln\left(\frac{1}{n}\left(e^{\sigma^{2}}-1\right)+1\right);\quad\mu_{\Sigma}=\ln n+(\sigma^{2}-\sigma_{\Sigma}^{2})/2.$
@@ B.3 Proof of Corollary 4
(Length Collapse in Text Embeddings) Given two texts of length $n$, the cosine similarity of their text embeddings tends to increase as $n$ grows.

### v2

@@ 6 Related Work
Text Embedding Models. Embeddings are generated by pre-trained language models (PLMs), which produce fixed-dimensional embeddings regardless of the input text length, laying the foundation of numerous NLP applications. Early works on text embeddings lack context awareness and are thus commonly labeled as word embedding models (Pennington et al., 2014). Modern embedding models (Wang et al., 2022a; Xiao et al., 2023) incorporate context awareness into language models through self-attention mechanisms, serving as the foundation for the latest embedding models. These models, based on Transformer architectures like BERT (Kenton and Toutanova, 2019) and RoBERTa (Liu et al., 2019), are first pre-trained on large-scale weakly supervised text pairs using contrastive loss (Gao et al., 2021), then fine-tuned on small scale but high-quality datasets. More recently, (Muennighoff et al., ) investigates the integration of generative and embedding tasks in large language models, introducing transformer-based GritLM, which enhances performance in both areas. Among the diverse range of pre-trained transformers (Wolf et al., 2020), self-attention plays a crucial role, and this paper focuses on how this module contributes to length collapse.
@@
Context Window Extension for Embedding Models. Despite transformer-based embedding models excelling in generating vector representations, they are typically constrained by a narrow context window of around 512 input tokens (Wang et al., 2022a; Xiao et al., 2023; Ni et al., 2022). This limitation significantly restricts their use in scenarios that require processing long inputs, such as extensive Wikipedia entries or meeting transcripts (Saad-Falcon et al., 2024; Zhu et al., 2024). Previous work attributes this poor performance to limited context windows and attempts to extend the window size. Current initiatives to develop long-context embedding models generally begin with acquiring a long-context backbone model, either by pre-training from scratch with long inputs (Günther et al., 2023; Nussbaum et al., 2024; Chen et al., 2024a) or utilizing existing models (Wang et al., 2024; Zhu et al., 2024). This is followed by training the backbone model to generate embeddings. However, in this paper, we discover that regardless of the model’s context window size, the model consistently performs worse on longer texts than on shorter texts due to length collapse. Therefore, we aim to improve the performance of long texts across all context window size models by analyzing and addressing length collapse, rather than simply expanding context window size.
7 Discussion
LLM-based Embedding Models: While our main focus is on PLM-based models, we also observe similar signs of Length Collapse in LLM-based models, which represent another major class of current mainstream models. We provide a detailed discussion of these observations in Appendix D. Contrastive Learning: Length Collapse increases similarity between long and short text embeddings. Although contrastive learning mitigates high similarity, performance degradation in long texts is also due to distribution differences between long and short texts, not just similarity. Further analysis of why contrastive learning fails to fully mitigate Length Collapse is provided in Appendix E.1. Comparison with Other Methods: We compare our method with post-processing approaches (Flow Function Li et al. (2020), Whitening Su et al. (2021)) and long-text methods (Chiang and Cholak (2022), YaRN Peng et al. (2024a)) in Appendix E.2 and E.3. Despite similar forms, these methods fail to address Length Collapse effectively.
8 Conclusion
@@
In this paper, we identify the phenomenon of Length Collapse, where the embeddings of longer texts tend to cluster together, and provide a Fourier domain analysis to explain this behavior. Our theoretical findings suggest that Self-Attention inherently performs stronger low-pass filtering as the text length increases, leading to patch uniformity issues in longer sentences. Furthermore, the distributional differences between short and long text embeddings contribute to the performance decline of longer texts. To address this, we propose TempScale, which effectively balances the filtering rates for short and long texts. Our extensive experiments validate the accuracy of our analysis and the effectiveness of our method, significantly improving the performance of general embedding models on the MTEB and LongEmbed benchmarks.
Limitations
@@
Limitations include: 1) LLM-based embedding model: Although we have observed a Length Collapse in LLM-based embedding models, further analysis is needed to investigate how unidirectional attention mechanisms specifically contribute to this phenomenon. Additionally, LLMs also exhibit a performance drop on long texts during text generation, which can be seen as Length Collapse. However, in this work, we only focus on PLM-based models across various tasks in MTEB. Future work could attempt to explain the reasons behind the performance decline of LLMs on long text generation from the perspective of low-pass filtering; 2) Tuning method: The work in this paper relies on existing models and pre-trained parameters without using a training dataset. In future work, we will focus on tuning the temperature for additional improvements; 3) Analysis on more modules: We primarily investigate the impact of the self-attention module on Length Collapse in this paper. Moving forward, we plan to explore the effects of additional modules in transformers such as LayerNorm and FFN.
Acknowledgements
This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (62472426,62376275). Supported by fund for building world-class universities (disciplines) of Renmin University of China. Work partially done at Beijing Key Laboratory of Research on Large Models and Intelligent Governance, and Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.
@@ Appendix A Background Information about Fourier Analysis
Hereby, we can independently operate $\mathcal{DC}\left[{\cdot}\right]$ and $\mathcal{HC}\left[{\cdot}\right]$ on the echo channel using the matrices in Eqn. A. Then we can write $\mathcal{DC}\left[{\cdot}\right]$ as below:
@@ B.1 Proof of Theorem 2
Denote $\bm{B}=\begin{pmatrix}\bm{b_{1}}&\cdots&\bm{b_{n}}\end{pmatrix}$ and we have $\bm{AB}=\begin{pmatrix}\bm{Ab_{1}}&\cdots&\bm{Ab_{n}}\end{pmatrix}$. From the definition of the spectral norm, we have:$\lVert\bm{A}\rVert_{2}\geq\frac{\lVert\bm{Ab_{i}}\rVert_{2}}{\lVert\bm{b_{i}}\rVert_{2}}.$ Taking the average of the right-hand side, we obtain: $\lVert\bm{A}\rVert_{2}^{2}\geq\sum_{i=1}^{n}\frac{\lVert\bm{b_{i}}\rVert_{2}^{2}}{\lVert\bm{B}\rVert_{F}^{2}}\frac{\lVert\bm{Ab_{i}}\rVert_{2}^{2}}{\lVert\bm{b_{i}}\rVert_{2}^{2}}.$ This implies: $\lVert\bm{A}\rVert_{2}^{2}\lVert\bm{B}\rVert_{F}^{2}\geq\sum_{i=1}^{n}\lVert\bm{Ab_{i}}\rVert_{2}^{2}=\lVert\bm{AB}\rVert_{F}^{2}.$ Finally, the last step utilizes the result $\lVert\bm{A}\rVert^{2}_{F}=\sum_{i,j}|a_{ij}|^{2}=\sum_{j=1}^{n}\lVert\bm{a_{j}}\rVert^{2}_{2}$. Because both the spectral norm and the Frobenius norm of a matrix remain unchanged under transposition, we have $\lVert\bm{AB}\rVert_{F}=\lVert\bm{B^{\top}A^{\top}}\rVert_{F}\leq\lVert\bm{B^{\top}}\rVert_{2}\lVert\bm{A^{\top}}\rVert_{F}=\lVert\bm{A}\rVert_{F}\lVert\bm{B}\rVert_{2}$.
@@
$=\frac{1}{n}\left(\bm{I}-\frac{1}{n}\bm{11}^{T}\right)\bm{A11}^{\top}\bm{X}\bm{W}_{V}$ (9)
$+\left(\bm{I}-\frac{1}{n}\bm{11}^{T}\right)\bm{AHW}_{V}$ (10)
@@ B.2 Proof of Theorem 3
(Filter Rate of Different Input Length $n$) Let $\bm{XW}_{Q}$ and $\bm{XW}_{K}$ be a Gaussian matrix, where elements $q_{ij}\sim\mathcal{N}(0,\sigma_{q}^{2})$ and $k_{ij}\sim\mathcal{N}(0,\sigma_{k}^{2}),\forall i,j.$ Let $x_{ij}=\bm{q}_{i}^{\top}\bm{k}_{j}/\sqrt{d}$ the attention score of pair $i,j$, whose variance can be expressed as $\sigma_{s}^{2}=\sigma_{q}^{2}\sigma_{k}^{2}+C_{cross}$, where $C_{cross}$ is the cross-covariance of the squared queries and keys Goodman (1960). Then we have
@@
The Eqn. 18 leverages $\left\lVert\bm{I}-\frac{1}{n}\bm{11}^{T}\right\rVert_{2}=1$ and $\lVert\bm{A}\rVert_{2}\leq\lVert\bm{A}\rVert_{F}$. Now we need to upper bound $\lVert\bm{A}\rVert_{F}$. Generally, the product of two independent Gaussian variables has a density in the form of a modified Bessel function of the second kind Nahshan et al. (2024). When the vector dimensions are sufficiently large, the Central Limit Theorem implies that the distribution of the dot product between $\bm{q}_{i}$ and $\bm{k}_{j}$ can be approximated by a Gaussian distribution with zero mean and variance $\sigma^{2}_{s}$. As mentioned in Theorem 3, the variance of $\bm{q}_{i}^{\top}\bm{k}_{j}$ can be expressed as $\sigma^{2}=\sigma^{2}_{q}\sigma^{2}_{k}+C_{\rm cross}$, where $C_{\rm cross}=\text{Cov}(\bm{q}^{2},\bm{k}^{2})-\text{Cov}(\bm{q},\bm{k})^{2}$ is the cross-covariance of the squared queries and keys Goodman (1960). Thus we can suppose that each element $p_{j}\sim\mathcal{N}(0,\sigma_{s})$ in the matrix $\bm{XW_{Q}W_{K}^{\top}X^{T}}$ is independent, where $j\in(1,\cdots,n)$:
@@
$\sigma_{\Sigma}^{2}=\ln\left(\frac{1}{n}\left(e^{\sigma^{2}}-1\right)+1\right);$
$\quad\mu_{\Sigma}=\ln n+(\sigma^{2}-\sigma_{\Sigma}^{2})/2.$
@@ B.3 Proof of Corollary 4
(Length Collapse in Text Embeddings) Let $x_{1},x_{2}\in\mathbb{R}^{d}$ be text embeddings of two documents with length $n$, decomposed as:
$x_{i}=DC{x_{i}}+HC{x_{i}}$
Assume $\|HC{x_{i}}\|_{2}\leq\beta(n)$ where $\beta(n)=O(1/\sqrt{n})$, and $\|DC{x_{i}}\|_{2}\geq\alpha>0$ is a constant. Then the cosine similarity satisfies:
$\lim_{n\to\infty}\cos(x_{1},x_{2})=1$